OpenStudy (anonymous):

solve by elimination please show your work 7x+2y=10 -7x+y=-16

OpenStudy (anonymous):

7x + 2y = 10 -7x + y = - 16 -------------add 0 + 3y = -6 3y = - 6 y = - 2 now sub -2 in for y in either of the original equations 7x + 2y = 10 7x + 2(-2) = 10 7x - 4 = 10 7x = 10 + 4 7x = 14 x = 2 check... -7x + y = - 16 -7(2) + (-2) = - 16 -14 - 2 = - 16 -16 = -16 (correct) ANSWER : x = 2 and y = - 2......(2,-2)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome :)